OpenStudy (anonymous):

true or false: if log (x+4)=2, then e^2=x+4

OpenStudy (anonymous):

false

OpenStudy (anonymous):

because it is 10^2 = x+4

OpenStudy (saifoo.khan):

the reason of being false is: it is log with base 10. but on right side it's log with base e.

OpenStudy (anonymous):

It depends. Conventionally, log written on its own has base 10, but I've had lecturers who meant it as having base e; in which case it would be true.

OpenStudy (anonymous):

False. For that to be true, the given should be \[\log_{e} (x+4) = 2\] and \[e^2 = (x+4)\]